# Knowledge Sharing in Manufacturing using Large Language Models: User Evaluation and Model Benchmarking

### 10 Jan 2024 v1

@@ Abstract.
Managing knowledge efficiently is crucial for organizational success. In manufacturing, operating factories has become increasing knowledge-intensive putting strain on the factory’s capacity to train and support new operators. In this paper, we introduce a Large Language Model (LLM)-based system designed to use the extensive knowledge contained in factory documentation. The system aims to efficiently answer queries from operators and facilitate the sharing of new knowledge. To assess its effectiveness, we conducted an evaluation in a factory setting. The results of this evaluation demonstrated the system’s benefits; namely, in enabling quicker information retrieval and more efficient resolution of issues. However, the study also highlighted a preference for learning from a human expert when such an option is available. Furthermore, we benchmarked several closed and open-sourced LLMs for this system. GPT-4 consistently outperformed its counterparts, with open-source models like StableBeluga2 trailing closely, presenting an attractive option given its data privacy and customization benefits. Overall, this work offers preliminary insights for factories considering using LLM-tools for knowledge management.
@@ 1. Introduction
Managing knowledge has become invaluable in today’s manufacturing industry, and technology plays an ever-increasing role. Factories already generated a wealth of knowledge in the form of issue reports and machine documentation; however, up till now, these sources have been challenging to harness using AI. With the advent of Large Language Models, such as GPT-4, the potential to use technology to support operators in knowledge-intensive tasks has leaped forward. However, using LLMs comes with many (unknown) practical and socio-technical risks and challenges. Furthermore, a generic LLM has limited use in a factory, where specific knowledge is needed. As such, an LLM-powered tool for manufacturing would require fine-tuning the model or including context information in its prompt (e.g., sections from a machine manual). We built a fully functional LLM-powered tool for this work that finds relevant sections from uploaded documents and issue reports to answer operator queries. We conducted a user study at a factory to evaluate the usability and potential of the system. Furthermore, we benchmarked several LLMs for the system, including closed and open source options. LLMs are typically benchmarked by the factuality of their response to general knowledge questions, resulting in LLM leaderboards111https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard. However, the ability to answer general knowledge questions is less relevant to a manufacturing environment where very specific, dynamic knowledge is required. Our findings could help decide how factories use this technology in the future.
1.1. Industry 5.0
Industry 5.0, the latest phase of industrial development, places human beings at the forefront of manufacturing processes, emphasizing their skills, creativity, and problem-solving abilities (Alves et al., 2023). This human-centered approach addresses the dehumanization of work, which has been a challenge in previous industrial revolutions. Industry 5.0 seeks to create a more fulfilling and productive work environment by empowering the workforce.
One of the fundamental problems that Industry 5.0 addresses is the alienation and dissatisfaction experienced by workers as a result of automation and machine dominance (Alves et al., 2023). The rapid advancement of technology has often led to a perception that humans are mere operators, overshadowed by the efficiency and precision of machines. However, in Industry 5.0, humans regain their significance in manufacturing. By integrating advanced technologies such as robotics and artificial intelligence, Industry 5.0 enables humans to collaborate with machines, leveraging their complementary strengths. This collaboration enhances productivity and efficiency and acknowledges the value of human contributions.
@@
Human-centered manufacturing in Industry 5.0 focuses on providing a work environment that nurtures individuals’ creativity and problem-solving capabilities (Maddikunta et al., 2022). It encourages workers to think critically, innovate, and continuously learn. With machines handling repetitive and mundane tasks, human workers can dedicate their time and energy to more complex and intellectually stimulating activities. This shift could enhance job satisfaction and promote personal and professional growth, as workers could acquire new skills and engage in higher-level decision-making (Alves et al., 2023), (Xu et al., 2021). Emphasis on human-machine collaboration and the continuous emergence and refinement of technology increase the need for adequate human-computer interaction (Brückner et al., 2023). One of the approaches to address this topic is using conversational AI to assist humans in manufacturing (Wellsandt et al., 2021).
1.2. Large-Language Model-Powered Tools for Knowledge-Intensive Scenarios
@@
The pre-training of LLMs on large-scale, mixed-source corpora allows them to capture extensive knowledge (Zhao et al., 2023), and demonstrate proficiency in interpreting complex information (Jawahar et al., 2019), generating general reasoning and insights (Wei et al., 2022), and aiding knowledge-intensive decision-making in various contexts. Consequently, researchers have been exploring LLMs adoption in both general and domain-specific tasks.
@@
However, despite their advantages, the responses generated by LLMs in the mentioned scenarios may encounter two potential issues: (1) outdated information originating from the model’s pre-training date, and (2) inaccuracies in factual representation, known as “hallucination”(Bang et al., 2023; Zhao et al., 2023). To address these challenges and leverage the capabilities of LLMs in domain-specific knowledge-intensive tasks, chain-of-thought and few-shot prompting (Brown et al., 2020; Gao et al., 2021) techniques are utilized.
This section explores LLM-based tools designed to acquire information and support decision making in both non-manufacturing and manufacturing settings.
@@
Using few-shot prompting to retrieve information across diverse topics, (Semnani et al., 2023) introduced an open-domain LLM-powered chatbot called WikiChat. WikiChat utilizes a 7-stage pipeline of few-shot prompted LLM that suggests facts verified against Wikipedia, retrieves additional up-to-date information, and generates coherent responses. They used a hybrid human-and-LLM method to evaluate the chatbot on different topics for factuality, alignment with real-worth truths and verifiable facts, and conversationality, a compound metric that scores how informational, natural, non-repetitive, and temporally correct the response is. Their solution significantly outperforms GPT-3.5 in factuality, with an average improvement of 24.4% while staying on par in conversationality. Others have explored the capabilities of LLMs in domain-specific tasks such as extracting structured data from unstructured healthcare texts (Tang et al., 2023), providing medical advice (Nov et al., 2023), simplifying radiology reports (Jeblick et al., 2022), Legal Judgement Prediction from multilingual legal documents (Trautmann et al., 2022), and scientific writing (Alkaissi and McFarlane, 2023).
Several manufacturers are cautiously adopting LLMs, while seeking solutions to mitigate its associated risks. For example, (Mercedes-Benz, 2023) is utilizing AI with ChatGPT integrated through Azure OpenAI Service to enhance error identification and analysis, quality management, and process optimization in intelligent vehicle production. This AI-driven approach simplifies complex evaluations and presentations of production-related data and quality management for quality engineers through dialogue-based queries.
@@
Also, in manufacturing, researchers have explored LLM-powered tools in knowledge-intensive scenarios. (Xia et al., 2023) demonstrated how using in-context learning and injecting task-specific knowledge into an LLM can streamline intelligent planning and control of production processes. (Kernan Freire et al., 2023) built a proof of concept for bridging knowledge gaps among workers by utilizing domain-specific texts, user requests, and query results from a knowledge graph and investigated the opportunities, risks, and user acceptance in two factories. (Wang et al., 2023) conducted a systematic test of ChatGPT’s responses to 100 questions sourced from course materials and industrial documents. They used zero-shot method and examined the responses’ correctness, relevance, clarity, and comparability. Their results suggested areas for improvement, including low scores when responding to critical analysis questions, occasional non-factual or out-of-manufacturing scope responses, and dependency on the query quality. Although (Wang et al., 2023) provides a comprehensive review of ChatGPT’s abilities to answer questions related to manufacturing, it did not include the injection of task-specific knowledge into the prompts.
@@
1.3. Evaluating Large-Language Models
@@
Large Language Model evaluation requires the definition of evaluation criteria, metrics, and datasets associated with the system’s main tasks. There are two types of LLM evaluations: intrinsic and extrinsic evaluation. Intrinsic evaluation focuses on the internal properties of a Language Model (Wei et al., 2023). It means the patterns and language structures learned during the pre-training phase. Extrinsic evaluation focuses on the model’s performance in downstream tasks, i.e., in the execution of specific tasks that make use of the linguistic knowledge gained upstream, like code completion (Xu et al., 2022). Despite extrinsic evaluation being computationally expensive, only conducting intrinsic evaluation is not comprehensive, as it only tests the LLMs capability for memorization (Jang et al., 2022). We will focus on extrinsic evaluation as we are primarily interested in the performance of LLM-based tools for specific real-world tasks,
@@
Extrinsic evaluation implies assessing the systems’s performance in tasks such as question answering, translation, reading comprehension, and text classification, among others (Kwon and Mihindukulasooriya, 2022). Existing benchmarks such as LAMBADA, HellaSwag, TriviaQA, BLOOM, Galactica, ClariQ and MMLU, among others, are widely reported in the literature for comparing language models. Likewise, domain-specific Benchmarks for tasks such as medical (Singhal et al., 2022), fairness evaluation (Zhang et al., 2023), finance (Xie et al., 2023), robot policies (Liang et al., 2022), and 3D printing code generation (Badini et al., 2023) can also be found.
Experts also evaluate Large-Language Models’ performance in specific downstream tasks. For example, (Lahat et al., 2023) evaluated the potential of ChatGPT for research priority identification. An expert panel was formed to review and rate the research questions generated by the system. Relevance, clarity, specificity, and originality were evaluated through a Likert scale. The interclass correlation coefficient (ICC) was also implemented to assess the inter-rater reliability.
@@
Numerous benchmarks exist; as outlined above, they range from specific downstream tasks to general language tasks. However, to our knowledge, LLM models have not been benchmarked for answering questions in the manufacturing domain based on information retrieved from provided context material, a technique known as Retrieval Augmented Generation (Lewis et al., 2020). Material such as machine documentation, standard work instructions, or issue reports will contain domain jargon and technical information that LLMs may struggle to process. Furthermore, the text in an issue report may pose additional challenges due to poor grammar or abbreviations. Therefore, as part of this work, we benchmarked several LLM models on their ability to answer questions based on factory manuals and issue reports. Furthermore, we conducted preliminary evaluations with factory managers to assess the potential benefits, risks and challenges. The following research questions informed our study:
RQ1
@@
What are the perceived benefits, challenges and risks of using Retrieval Augmented Generation for information retrieval and knowledge sharing for factory operators?
RQ2
@@
How do Large Language Models (e.g., GPT-4, GPT-3.5, Llama2) compare in performance when employing Retrieval Augmented Generation over factory documentation to answer queries from operators? We consider performance as the factuality, completeness, hallucinations, and conciseness of the generated response.
@@ 2. System
We built a fully functional system to assess the potential of using LLMs for information retrieval and knowledge sharing for factory operators. Benefiting from LLMs’ in-context learning capabilities, we use this to supply an LLM with information in the form of factory manuals, and issue reports relevant to the user’s question, a technique known as Retrieval Augmented Generation (RAG), see Figure 1. As pointed out by (Wei et al., 2022), training LLMs using a prompt packed with query-related information can yield substantial performance enhancement. Users can ask questions in the chat box by typing or using voice input. The response is displayed at the top of the page, and the document chunks used for the answer can be checked at the bottom (see Figure 2).
2.0.1. Tool Dependencies
@@
2.0.2. Knowledge Base Construction
Our experiment incorporates two distinct types of domain-specific data: factory manuals and shared knowledge from factory workers. Factory manuals outline information on machine operation, safety protocols, quality assurance, and more. These resources, provided by factory management teams, initialize the knowledge base for each specific factory. The materials come in various formats, including PDF, Word, and CSV files. To ensure easy access and indexing, we leverage the capabilities of LlamaIndex.
In addition to the manuals, we integrate valuable practical knowledge from factory workers. This information is gathered from the production line, utilizing the five-why process, an iterative technique cited by (Serrat, 2017). The five-why technique probes into cause-and-effect relationships underlying specific problems by repeatedly asking ”Why?” until the root cause is revealed, typically by the fifth query. This process enables us to gather real-world issues encountered on production lines, which may not be covered in the factory manuals.
Since these documents can be extensive, integrating them directly into the LLM conversation could be challenging. To overcome this, we further segment the materials into manageable chunks, each comprising approximately 400 tokens. This method effectively incorporates the materials into the LLM prompt without compromising the conversation flow. Following the segmentation, each document chunk is processed through LlamaIndex using the OpenAI Embedding API222https://api.openai.com/v1/embeddings. Utilizing the ‘text-embedding-ada-002’ model, LlamaIndex transforms each chunk into a corresponding embedding vector. These resulting vectors are then securely stored, ready for future retrieval and use.
2.0.3. Query Construction
To retrieve the document data relevant to specific user queries, we employ the same embedding model, “text-embedding-ada-002”, to generate vector representations of these queries. By leveraging the similarity calculation algorithm provided by LlamaIndex, we can identify and retrieve the top-K most similar segmented document snippets related to the user query. This allows us to construct pertinent LLM queries. Once the snippets are retrieved, they are synthesized into the following query template based on the templates used by LlamaIndex333https://docs.llamaindex.ai:

### 26 Feb 2024 v2

@@ Abstract.
Recent advances in natural language processing enable more intelligent ways to support knowledge sharing in factories. In manufacturing, operating production lines has become increasingly knowledge-intensive, putting strain on a factory’s capacity to train and support new operators. This paper introduces a Large Language Model (LLM)-based system designed to retrieve information from the extensive knowledge contained in factory documentation and knowledge shared by expert operators. The system aims to efficiently answer queries from operators and facilitate the sharing of new knowledge. We conducted a user study at a factory to assess its potential impact and adoption, eliciting several perceived benefits, namely, enabling quicker information retrieval and more efficient resolution of issues. However, the study also highlighted a preference for learning from a human expert when such an option is available. Furthermore, we benchmarked several commercial and open-sourced LLMs for this system. The current state-of-the-art model, GPT-4, consistently outperformed its counterparts, with open-source models trailing closely, presenting an attractive option given their data privacy and customization benefits. In summary, this work offers preliminary insights and a system design for factories considering using LLM tools for knowledge management.
@@ 1. Introduction
Human-centric manufacturing seeks to support human operators with technology, aiming to enhance creativity, human well-being, problem-solving abilities, and overall productivity within factories (Alves et al., 2023; May et al., 2015; Fantini et al., 2020). However, a significant challenge persists in effectively managing and utilizing the vast knowledge generated within these manufacturing environments, such as issue reports and machine documentation (Man, 2014). This knowledge is crucial for optimizing operations, yet it remains largely untapped due to the difficulties in processing and interpreting the disconnected, sometimes unstructured, technical information it contains (Leoni et al., 2022; Edwards et al., 2008).
Traditionally, leveraging this knowledge has been cumbersome, with operators choosing to use personal smartphones over official procedures (Richter et al., 2019) and AI unable to handle the complexity of the data (Edwards et al., 2008). However, recent Large Language Models (LLMs) like GPT-4 show promise in addressing these challenges. LLMs can effectively interpret, summarize, and retrieve information from vast text-based datasets (Lewis et al., 2020) while concurrently aiding the capture of new knowledge (Kernan Freire et al., 2023b). These capabilities could significantly support operators in knowledge-intensive tasks, making it easier to access relevant information, share new knowledge, and make informed decisions rapidly.
While LLMs offer promising capabilities, their application in manufacturing is not straightforward. The specific, dynamic knowledge required in this domain poses unique challenges (Feng et al., 2017). For instance, a foundational LLM may have limited utility in a factory setting without significant customization, such as fine-tuning or incorporating specific context information into its prompts (Wang et al., 2023b). Additionally, the practical and socio-technical risks and challenges of deploying LLMs in such environments remain largely unexplored — factors key to human-centered AI (Shneiderman, 2022). Concerns include the accuracy of the information provided, the potential for ”hallucinated” answers (Zuccon et al., 2023), and the need for systems that can adapt to the highly specialized and evolving knowledge base of a specific manufacturing setting (Feng et al., 2017).
In response to these challenges, we developed an LLM-powered tool to leverage factory documents and issue analysis reports to answer operators’ queries. Furthermore, the tool facilitates the analysis and reporting of new issues. This tool demonstrates the feasibility of using LLMs to enhance knowledge management in manufacturing settings. To understand its effectiveness and potential, we conducted a user study in a factory environment, evaluating the system’s usability, user perceptions, adoption, and impact on factory operations.
Our approach also addresses the lack of specific benchmarks for evaluating LLMs in manufacturing. We benchmarked several LLMs, including both closed and open-source models, recognizing that the standard benchmarks111https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard—last accessed February 26, 2024 primarily focus on general knowledge and reasoning. As such, they may not adequately reflect the challenges of understanding manufacturing-specific terminology and concepts. This benchmarking focused on their ability to utilize factory-specific documents and unstructured issue reports to provide factual and complete answers to operators’ queries.
2. Background
In this section, we address the topic of industry 5.0, LLM-powered tools for knowledge management, benchmarking LLMs, and the research questions informing this work.
2.1. Human-centered Manufacturing
@@
Industry 5.0, the latest phase of industrial development, places human beings at the forefront of manufacturing processes, emphasizing their skills, creativity, and problem-solving abilities (Xu et al., 2021a; Maddikunta et al., 2022b; Alves et al., 2023). Human-centered manufacturing in Industry 5.0 focuses on providing a work environment that nurtures individuals’ creativity and problem-solving capabilities (Maddikunta et al., 2022a). It encourages workers to think critically, innovate, and continuously learn. With machines handling repetitive and mundane tasks, human workers can dedicate their time and energy to more complex and intellectually stimulating activities. This shift could enhance job satisfaction and promote personal and professional growth, as workers could acquire new skills and engage in higher-level decision-making (Alves et al., 2023; Xu et al., 2021b). Emphasis on human-machine collaboration and the continuous emergence and refinement of technology increases the need for adequate human-computer interaction (Brückner et al., 2023). One of the approaches to address this topic is using conversational AI to assist humans in manufacturing (Wellsandt et al., 2021).
2.2. LLM-powered Knowledge Management Tools
@@
Training Large Language Models (LLMs) on numerous, diverse texts results in the embedding of extensive knowledge (Zhao et al., 2023). LLMs can also adeptly interpret complex information (Jawahar et al., 2019), general reasoning (Wei et al., 2022a), and aiding knowledge-intensive decision-making. Consequently, researchers have been exploring applying LLM-powered tools in domain-specific tasks (Xie et al., 2023b; Wen et al., 2023; Zhang et al., 2023b).
@@
Despite their potential benefits, the responses generated by LLMs may have two potential issues: (1) outdated information originating from the model’s training date, and (2) inaccuracies in factual representation, known as “hallucinations” (Bang et al., 2023; Zhao et al., 2023). To address these challenges and leverage the capabilities of LLMs in domain-specific knowledge-intensive tasks, several techniques can be used, such as chain-of-thought (Wei et al., 2022b), few-shot prompting (Brown et al., 2020; Gao et al., 2021), and retrieval augmented generation (Lewis et al., 2020).
@@
Using few-shot prompting to retrieve information across diverse topics, Semnani et al. (2023) introduced an open-domain LLM-powered chatbot called WikiChat. WikiChat utilizes a 7-stage pipeline of few-shot prompted LLM that suggests facts verified against Wikipedia, retrieves additional up-to-date information, and generates coherent responses. They used a hybrid human-and-LLM method to evaluate the chatbot on different topics for factuality, alignment with real-worth truths and verifiable facts, and conversationality. This compound metric scores how informational, natural, non-repetitive, and temporally correct the response is. Their solution significantly outperforms GPT-3.5 in factuality, with an average improvement of 24.4% while staying on par in conversationality. Others have explored the capabilities of LLMs in domain-specific tasks such as extracting structured data from unstructured healthcare texts (Tang et al., 2023), providing medical advice (Nov et al., 2023), simplifying radiology reports (Jeblick et al., 2022), Legal Judgement Prediction from multilingual legal documents (Trautmann et al., 2022), and scientific writing (Alkaissi and McFarlane, 2023).
@@
Several manufacturers are cautiously adopting LLMs, while seeking solutions to mitigate their associated risks. For example, Mercedes-Benz (2023) used AI with ChatGPT integrated through Azure OpenAI Service to enhance quality management and process optimization in vehicle production. This AI-driven approach simplifies complex evaluations for quality engineers through dialogue-based queries. Xia et al. (2023) demonstrated how using in-context learning and injecting task-specific knowledge into an LLM can streamline intelligent planning and control of production processes. Kernan Freire et al. (2023a) built a proof of concept for bridging knowledge gaps among workers by utilizing domain-specific texts and knowledge graphs. Wang et al. (2023a) conducted a systematic test of ChatGPT’s responses to 100 questions from course materials and industrial documents. They used a zero-shot method and examined the responses’ correctness, relevance, clarity, and comparability. Their results suggested areas for improvement, including low scores when responding to critical analysis questions, occasional non-factual or out-of-manufacturing scope responses, and dependency on query quality. Although (Wang et al., 2023a) provides a comprehensive review of ChatGPT’s abilities to answer questions related to manufacturing; it did not include the injection of task-specific knowledge into the prompts.
@@
2.3. Evaluating LLMs
@@
Large Language Model evaluation requires the definition of evaluation criteria, metrics, and datasets associated with the system’s main tasks. There are two types of LLM evaluations: intrinsic and extrinsic evaluation. Intrinsic evaluation focuses on the internal properties of a Language Model (Wei et al., 2023). It means the patterns and language structures learned during the pre-training phase. Extrinsic evaluation focuses on the model’s performance in downstream tasks, i.e., in the execution of specific tasks that make use of the linguistic knowledge gained upstream, like code completion (Xu et al., 2022). Despite extrinsic evaluation being computationally expensive, only conducting intrinsic evaluation is not comprehensive, as it only tests the LLMs capability for memorization (Jang et al., 2022). Here, we focus on extrinsic evaluation as we are primarily interested in the performance of LLM-based tools for specific real-world tasks.
@@
Extrinsic evaluation implies assessing the systems’s performance in tasks such as question answering, translation, reading comprehension, and text classification, among others (Kwon and Mihindukulasooriya, 2022). Existing benchmarks such as LAMBADA, HellaSwag, TriviaQA, BLOOM, Galactica, ClariQ and MMLU, among others, are widely reported in the literature for comparing language models. Likewise, domain-specific Benchmarks for tasks such as medical (Singhal et al., 2022), fairness evaluation (Zhang et al., 2023a), finance (Xie et al., 2023a), robot policies (Liang et al., 2022), and 3D printing code generation (Badini et al., 2023) can also be found. Experts also evaluate the performance of large-language models (LLMs) in specific downstream tasks, such as using physicians to evaluate the output of medical specific LLMs (Singhal et al., 2022).
@@
LLM benchmarks range from specific downstream tasks to general language tasks. However, to our knowledge, LLMs have not been benchmarked for answering questions in the manufacturing domain based on context material, a technique known as Retrieval Augmented Generation (Lewis et al., 2020). Material such as machine documentation, standard work instructions, or issue reports will contain domain jargon and technical information that LLMs may struggle to process. Furthermore, the text in an issue report may pose additional challenges due to abbreviations, poor grammar, and formatting (Edwards et al., 2008). Therefore, as part of this work, we benchmarked several LLMs on their ability to answer questions based on factory manuals and unstructured issue reports. Furthermore, we conducted a user study with factory operators and managers to assess the potential benefits, risks and challenges. The following research questions informed our study:
@@
What are the perceived benefits, challenges, and risks of using Large Language Models for information retrieval and knowledge sharing for factory operators?
@@
How do Large Language Models compare in performance when answering factory operators’ queries based on factory documentation and unstructured issue reports? We consider performance as the factuality, completeness, hallucinations, and conciseness of the generated response.
@@ 3. System
We built a fully functional system to assess the potential of using LLMs for information retrieval and knowledge sharing for factory operators. Benefiting from LLMs’ in-context learning capabilities, we use this to supply an LLM with information in the form of factory manuals, and issue reports relevant to the user’s question, a technique known as Retrieval Augmented Generation (RAG) (Lewis et al., 2020), see Figure 1. As noted by Wei et al. (2022a), training LLMs using a prompt packed with query-related information can yield substantial performance enhancement. Users can ask questions in the chat box by typing or using voice input. The response is displayed at the top of the page, and the document chunks used for the answer can be checked at the bottom (see Figure 2).
3.1. Tool Dependencies
@@
Since the factory documents can be long, they may overflow the LLM’s context window or result in unnecessary computational demand. To overcome this, we segment the materials into manageable chunks, each comprising approximately 400 tokens. This method effectively incorporates the materials into the LLM prompt without compromising the conversation flow. Following the segmentation, each document chunk is processed through LlamaIndex using the OpenAI Embedding API222https://api.openai.com/v1/embeddings—last accessed February 26, 2024. Utilizing the “text-embedding-ada-002” model, LlamaIndex transforms each chunk into a corresponding embedding vector. These resulting vectors are then securely stored, ready for future retrieval and use.
3.2. Knowledge Base Construction
Our experiment incorporates two distinct types of domain-specific data: factory manuals and shared knowledge from factory workers. Factory manuals outline information on machine operation, safety protocols, quality assurance, and more. These resources, provided by factory management teams, initialize the knowledge base for each specific factory. The materials come in various formats, including PDF, Word, and CSV files.
In addition to the factory manuals, we integrate issue analysis reports from factory workers. This information is gathered from the production line, utilizing the five-why process, an iterative root-cause analysis technique (Serrat, 2017) (right side of Figure 2). The five-why technique probes into cause-and-effect relationships underlying specific problems by repeatedly asking ”Why?” until the root cause is revealed, typically by the fifth query. This process enables us to gather real-world issues encountered on production lines, which may not be covered in the factory manuals. Upon entering all required information, including one or more “whys”, the operator presses “check”, triggering a prompt to the LLM that performs a logical check of the entered information and checks for inconsistencies with previously reported information. The operator can revise the entered information and submit it as is. Then, the submitted report will be added to a queue for expert operators to check before it is added to the knowledge base.
3.3. Query Construction
To retrieve the document data relevant to specific user queries, we employ the same embedding model, “text-embedding-ada-002” to generate vector representations of these queries. By leveraging the similarity calculation algorithm provided by LlamaIndex, we can identify and retrieve the top-K most similar segmented document snippets related to the user query. This allows us to construct pertinent LLM queries. Once the snippets are retrieved, they are synthesized into the following query template based on the templates used by LlamaIndex333https://docs.llamaindex.ai—last accessed February 26, 2024: